chazz5:

Use substitution to solve the system. 3x+5y=29 -2x+y=-2 x= y=

chazz5:

@arieonna

Gayass:

32583475832

chazz5:

thats not right

Gayass:

it is

Gayass:

how would you know

Gayass:

your the one asking

Gayass:

try it

Arieonna:

3x+5y=2,9-2x+y=y:x=9/2, y=23/10 hope this is what u need

chazz5:

thx @arieonna im taking a quiz i need help

Gayass:

its wrong

Arieonna:

u bet math can be chalenging i use to be really good at math until i came to remote learning

Gayass:

you need to flip with wy